Calculate the hydronium ion concentration and pH in a 0.92 M solution of ammonium chloride.

Ka NH4+ = 5.6x10^-10
Answer #1

NH4+ dissociates as:

NH4+ -----> H+ + NH3

0.92 0 0

0.92-x x x

Ka = [H+][NH3]/[NH4+]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((5.6*10^-10)*0.92) = 2.27*10^-5

since c is much greater than x, our assumption is correct

so, x = 2.27*10^-5 M

So, [H+] = x = 2.27*10^-5 M

use:

pH = -log [H+]

= -log (2.27*10^-5)

= 4.644

Answer:

[H+] = 2.27*10^-5 M

pH = 4.64